Hira has Rs.1200 a week to spend on yogurt and mangoes. The price of yogurt is Rs.200, and mangoes are Rs.400 a box. (Hint: Quantity of mangoes to be plotted on the x-axis of the budget line)


How do Hira’s budget line change if, other things remaining the same,


Rotate outward

Rotate inward

Shift parallel outward

Shift parallel outward

Bend inside from the middle

Bend outside from the middle

The price of mangoes falls?


Hira’s income increases?

         


      

The price of yogurt rises?

         


Hira’s income decreases?

A.Shift parallel outward

B.Shift parallel inward

C.Rotate outward

D.Rotate inward

E.Bend outside from the middle

F.Bend inside from the middle



1
Expert's answer
2021-06-02T12:05:48-0400

Solution:

The Y-intercept for the budget line ="\frac{M}{Py}"

M = Total income = Rs.1,200

Py = Price of yogurt = Rs.200


The Y-intercept = "\frac{1200}{200} = 6"

 

The X-intercept for the budget line = "\frac{M}{Px}"

M = Total income = Rs.1,200

Px = Price of mangoes = Rs.400


The X-intercept = "\frac{1200}{400} = 3"

M = PxX + PyY

1,200 = 400X + 200Y

 

 The changes are explained by the below graph:



 

 

 

 

a.). If the price of mangoes falls - D. Rotate outward

The budget line will rotate outward by shifting to the right to the new position BL1, since consumers purchase more of good x.

 

b.). Hira’s income increases - E. Shift parallel outward

The budget line will shift outward parallel to the original budget line to BL3 since consumers will be able to purchase more of good x and good y due to increased income levels.

 

c.). The price of yogurt rises - C. Rotate inward

The budget line will rotate inward by shifting to the left to the new position BL2 since consumers will purchase less of good y.


d.). Hira’s income decreases - B. Shift parallel inward

The budget line will shift inward parallel to the original line to BL4 since consumers can purchase less of both goods with less income.